A four stroke cycle diesel engine has a brake thermal efficiency of 0.32 and lower heat of combustion of fuel is 41200kJ/kg. Calculate the brake specific fuel of consumption and the brake mean effective pressure of the engine if the volumetric efficiency is 0.76, inlet density of charge is 1.05 kg/m3 and fuel air ratio


0.04,


is


1
Expert's answer
2021-03-08T03:09:30-0500

Brake thermal efficiency = 32%

Lower heat combustion = 41200kJ/kg

Volumetric efficiency = 76%

Density of charge = 1.05 kg/m³

Fuel air ratio = 0.04


Specific fuel consumption = ?

Brake mean effective pressure of engine = ?


Specific fuel combustion = 1/0.04 × 1/41200 × 76/100 = 5 × 10-⁴ kg/kw-hr = 0.5g/kw-hr


Brake mean effective pressure of engine = 76/100 × 1.05 × 41200 = 32877.6 Pa